Answer all, show all work:

Ervin Company uses the allowance method to account for uncollectible accounts receivable. Bad debt expense is established as a percentage of credit sales. For 2021, net credit sales totaled $5,600,000, and the estimated bad debt percentage is 1.70%. No previously written-off accounts receivable were reinstated during 2021. The allowance for uncollectible accounts had a credit balance of $53,000 at the beginning of 2021 and $45,500, after adjusting entries, at the end of 2021.

Required:
1.
What is bad debt expense for 2021 as a percent of net credit sales?
2. Assume Ervin makes no other adjustment of bad debt expense during 2021. Determine the amount of accounts receivable written off during 2021.
3. If the company uses the direct write-off method, what would bad debt expense be for 2021?
Answer #1

1) Bad debt expense = 5600000*1.7% = 95200

2) Amount write off = 53000+95200-45500 = 102700

3) Bad debt expense = 102700